OpenStudy (anonymous):

Write this equation in standard form. Then use the x-intercept and y-intercept to graph the equation. y=2x-10

OpenStudy (anonymous):

2x-y-10=0 And then set x=0 to find y when x=0 (y=10) And then set y=0 to find x when y=0 (x=5) And to graph just join a line between (0,10) and (5,0)